# Zero-shot Concept Bottleneck Models

### v1

@@ Abstract
Figure 1: Zero-shot concept bottleneck models (Z-CBMs). Z-CBMs predict concepts for input by retrieving them from a large-scale concept bank. Then, Z-CBMs predict labels based on the weighted sum of the retrieved concept vectors with importance weights yielded by sparse linear regression.
@@ 1 Introduction
One of the primary interests of the deep learning research community is developing a human-interpretable model. Concept bottleneck model (CBM, Koh et al. (2020)) is an inherently interpretable neural network model, which aims to explain their final prediction via the intermediate concept predictions. CBMs are trained on a target task to learn the input-to-concept and concept-to-label mappings in an end-to-end manner. A concept is composed of high-level semantic vocabulary for describing objects of interest in input data. For instance, CBMs can predict the final label “apple” from the linear combination of the concepts “red sphere,” ”green leaf,” and “glossy surface.” These intermediate concept predictions not only provide interpretability but also intervenability in the final prediction by editing the predicted concepts.
@@
In the original CBMs (Koh et al., 2020), a concept set for each class label is defined by manual annotations, incurring massive labeling costs greater than ones of the class labels. To reduce the costs, Oikarinen et al. (2023) and Yuksekgonul et al. (2023) automatically generate the concept sets by large language models (LLMs, e.g., GPT-3 (Brown et al., 2020a)) and use the multi-modal embedding space of vision-language models (VLMs, e.g., CLIP (Radford et al., 2021)) to learn the input-to-concept mapping through similarities in the multi-modal feature space. Although modern CBMs are free from manual pre-defined concepts, we argue that the practicality is still restricted by the requirements of training input-to-concept and concept-to-label mappings on target datasets. In other words, CBMs are not available without manually collecting target datasets and additional training of model parameters on them.
@@
To overcome this limitation, this paper tackles a new problem setting of CBMs in a zero-shot manner for target tasks. In this setting, we can access pre-trained VLMs, but we cannot know the concepts composing target data in advance. This setting forces models to perform two-stage zero-shot inference of input-to-concept and concept-to-label for unseen input samples. The zero-shot input-to-concept inference can not be solved by a naïve application of VLMs as the ordinary zero-shot classification of input-to-label because it is required to infer a subset of relevant concepts, not a single label, from the large set of all concepts. Furthermore, the zero-shot concept-to-label inference is difficult because the concept-to-label mapping is not obvious without target data and training, which are unavailable in this setting. Therefore, we aim to answer the following research question: how can we provide interpretable and intervenable concepts by the zero-shot input-to-concept/concept-to-label inference without target datasets and training?
@@
We present a novel CBM class called zero-shot concept bottleneck models (Z-CBMs). Z-CBMs are zero-shot interpretable models that employ off-the-shelf pre-trained VLMs with frozen weights as the backbone (Fig. 1). Conceptually, Z-CBMs dynamically find concepts related to the input from a broad concept bank (concept retrieval) and then predict the final label by simulating zero-shot classification of black-box VLMs via reconstructing the original input embedding from the concept embeddings (concept regression). Our primary contribution is to achieve zero-shot input-to-concept and concept-to-label inference with this framework without additional training.
@@
We implement the components of Z-CBMs with simple yet carefully designed and effective techniques. For concept retrieval, Z-CBMs should cover broad domains to provide sufficient concepts for unseen inputs. To cover broad concepts, we build a large-scale concept bank, which is composed of millions of vocabulary extracted from large-scale text caption datasets such as YFCC (Thomee et al., 2016). Given an input sample, Z-CBMs dynamically retrieve concept candidates from the concept bank with an efficient and scalable cross-modal search algorithm. For concept regression, Z-CBMs estimate the importance of concepts for the input feature and then predict labels by the importance-weighted concept features. However, many of the retrieved concept candidates semantically overlap each other, and thus, the semantically duplicated concepts with high importance by a naïve estimation method can harm the interpretability and intervenability for humans. To overcome this challenge, Z-CBMs find essential and mutually exclusive concepts for the final label prediction by leveraging sparse linear regression (e.g., lasso) to reconstruct the input visual feature vector by a weighted sum of the concept candidate vectors. Combining concept retrieval and concept regression enables Z-CBMs to predict final task labels with interpretable concepts for various domain inputs without any target task training.
@@
Our extensive experiments on 12 datasets demonstrate that Z-CBMs can provide interpretable and intervenable concepts without any additional training. Specifically, we confirm that Z-CBMs’ sparse concepts are well correlated to input images and cover the annotated concepts in the existing training-based CBMs. Furthermore, the performance of Z-CBMs can be enhanced by human intervention in the predicted concepts, emphasizing the reliability of the concept-based prediction. We also show that Z-CBMs can perform accurately at a competitive level with black box VLMs and existing CBMs with training. These results suggest the practicality of Z-CBMs for various domains.
2 Related Work
CBMs (Koh et al., 2020) are inherently interpretable deep neural network models that predict concept labels and then predict final class labels from the predicted concepts. In contrast to the other explanation styles such as post-hoc attribution heatmaps (Lundberg & Lee, 2017; Selvaraju et al., 2017; Sundararajan et al., 2017), CBMs provide semantic ingredients consisting the final label prediction through the bilevel prediction of input-to-concept and concept-to-label. The original CBMs have the challenge of requiring human annotations of concept labels, which are more difficult to obtain than target task labels. Another challenge is the performance degradation from backbone black-box models (Zarlenga et al., 2022; Moayeri et al., 2023; Xu et al., 2024) due to the difficulty of learning long-tailed concept distributions (Ramaswamy et al., 2023). Post-hoc CBMs (Yuksekgonul et al., 2023), Label-free CBMs (Oikarinen et al., 2023), and LaBo (Yang et al., 2023) addressed these challenges by automatically collecting concepts corresponding to target task labels by querying LLMs (e.g., GPT-3 (Brown et al., 2020b)) and leveraging multi-modal feature spaces of pre-trained VLMs (e.g., CLIP (Radford et al., 2021)) for learning the input-to-concept mapping. Subsequently, the successor works have basically assumed the use of LLMs or VLMs, further advancing CBMs (Panousis et al., 2023; Rao et al., 2024b; Tan et al., 2024; Srivastava et al., 2024). In particular, Panousis et al. (2023) and Rao et al. (2024a) are related to our work in terms of using sparse modeling to select concepts for input images. However, all of these existing CBMs still require training specialized neural networks on target datasets, incurring additional target data collection and training resources. Handling the bi-level prediction in a zero-shot manner for unseen input is unobvious because it can not be solved by naïve application of the existing zero-shot classification methods, which depend on task-specific vocabularies for single label predictions (Norouzi et al., 2014; Demirel et al., 2017; Menon & Vondrick, 2023). Furthermore, current CBMs and the recent interpretable framework for CLIP (Bhalla et al., 2024) limit the number of concepts up to a few thousand due to training and computational constraints, restricting the generality. In contrast to the previous CBMs, our Z-CBMs can perform fully zero-shot inference based on a large-scale concept bank with millions of vocabulary for arbitrary input images in various domains as shown in the experiments in Sec. 5.5.
@@ 3 Zero-shot Concept Bottleneck Models
In this section, we formalize the framework of Z-CBMs, which perform a zero-shot inference of input-to-concept and concept-to-label without target datasets and additional training (Fig. 1). Z-CBMs are composed of concept retrieval and concept regression. Concept retrieval finds a set of the most input-related concept candidates from millions of concepts by querying an input image feature with a semantic similarity search (Fig. 2). Concept regression estimates the importance scores of the concept candidates by sparse linear regression to reconstruct the input feature (Fig. 2). Finally, Z-CBMs provide the final label predicted by the reconstructed vector and concept explanations with importance scores.
3.1 Problem Setting
We inherit the problem setting of existing vision-language-based CBMs (Oikarinen et al., 2023) except for not updating any neural network parameters. The goal is to predict the final task label $y\in\mathcal{Y}$ of input $x\in\mathcal{X}$ based on $K$ interpretable textual concepts $\{c_{i}\in\mathcal{C}\subset\mathcal{T}\}_{i=1}^{K}$, where $\mathcal{X}$, $\mathcal{Y}$, $\mathcal{C}$, and $\mathcal{T}$ are the input, label, concept, and text space, respectively. To this end, we predict the final task label by the bi-level prediction $h\circ g(x)$, where $g:\mathcal{X}\to\mathcal{C}^{K}$ is a concept predictor and $h:\mathcal{C}^{K}\to\mathcal{Y}$ is a label predictor. This setting allows to access a vision encoder $f_{\mathrm{V}}:\mathcal{X}\to\mathbb{R}^{d}$ and a text encoder $f_{\mathrm{T}}:\mathcal{T}\to\mathbb{R}^{d}$ provided by a VLM like CLIP (Radford et al., 2021), and a concept bank $C=\{c_{i}\}^{N_{\mathrm{c}}}_{i=1}$. The concept bank $C$ is composed of unique concepts from arbitrary sources, including manually collected concepts and automatically generated concepts by LLMs like GPT-3 (Brown et al., 2020a).
@@ 3.1 Problem Setting
3.2 Zero-shot Inference
Concept Retrieval. We first find the most semantically closed concept candidates to input images from the large spaces in a concept bank (Fig. 2). Given an input $x$, we retrieve the set of $K$ concept candidates $C_{x}\subset C$ by using image and text encoders of pre-trained VLMs $f_{\mathrm{V}}$ and $f_{\mathrm{T}}$ as
$C_{x}=\underset{c\in C}{\operatorname{Ret}_{K}}(f_{\mathrm{V}}(x),f_{\mathrm{T}}(c))=\underset{c\in C}{\operatorname{Top-K}}~{}\operatorname{Sim}(f_{\mathrm{V}}(x),f_{\mathrm{T}}(c)),$ (1)
where ${\operatorname{Top-K}}$ is an operator yielding top-$K$ concepts in $C$ from a list sorted in descending order according to a similarity metric $\operatorname{Sim}$. Throughout this paper, we use cosine similarity as $\operatorname{Sim}$ by following (Conti et al., 2023). Thanks to the scalability of the similarity search algorithm (Johnson et al., 2019; Douze et al., 2024), Eq. (1) can efficiently find the concept candidates in an arbitrary concept bank $C$, which contains millions of concepts to describe inputs in various domains.
Concept Regression. Given a concept candidate set $C_{x}=\{c_{1},...,c_{K}\}$, we predict the final label $\hat{y}$ by selecting essential concepts from $C_{x}$. Conventional CBMs infer the mapping between $C_{x}$ and $\hat{y}$ by training neural regression parameters on target tasks, which incurs the requirements of target dataset collections and additional training costs. Instead, we solve this task with a different approach leveraging the zero-shot performance of VLMs. As shown in the previous studies (Radford et al., 2021; Jia et al., 2021), VLMs can be applied to zero-shot classification by inferring a label $\hat{y}$ by matching input $x$ and a class name text $t_{y}\in\mathcal{T}$ in the multi-modal feature spaces as follows.
@@ 3.2 Zero-shot Inference
Through this objective, we can achieve $W$ not only for approximating image features but also for effectively estimating the contribution of each concept to the label prediction owing to the sparse regularization $\|W\|_{1}$. Since $C_{x}$ is retrieved from large-scale concept bank $C$, it often contains noisy concepts that are similar to each other, undermining interpretability due to semantic duplication. In this sense, the sparse regularization enhances interpretability since it can eliminate unimportant concepts for the label prediction (Hastie et al., 2015).
@@
2:  $C_{x}\leftarrow\underset{c\in C}{\operatorname{Ret}_{K}}(f_{\mathrm{V}}(x),f_{\mathrm{T}}(c))$
3:  $F_{C_{x}}\leftarrow[f_{\mathrm{T}}(c_{1}),...,f_{\mathrm{T}}(c_{K})]$
@@ 4 Implementation
Vision-Language Models. Z-CBMs allow to leverage arbitrary pre-trained VLMs for $f_{\mathrm{V}}$ and $f_{\mathrm{T}}$. We basically use the official implementation of OpenAI CLIP (Radford et al., 2021) and the publicly available pre-trained weights. [^1] https://github.com/openai/CLIP Specifically, by default, we use ViT-B/32 as $f_{\mathrm{V}}$ and the base transformer with 63M parameters as $f_{\mathrm{T}}$ by following the original CLIP. In Section 5.6.1, we show that other VLM backbones (e.g., SigLIP (Zhai et al., 2023) and OpenCLIP (Cherti et al., 2023)) are also available for Z-CBMs.
Concept Bank Construction. Here, we introduce the construction protocols of the concept bank $C$ of Z-CBMs. Since Z-CBMs can not know concepts of input image features in advance, a concept bank should contain sufficient vocabulary to describe the various domain inputs. To this end, we extract concepts from multiple image caption datasets and integrate them into a single concept bank. Specifically, we automatically collect concepts as noun phrases by parsing each sentence in the caption datasets including Flickr-30K (Young et al., 2014), CC-3M (Sharma et al., 2018), CC-12M (Changpinyo et al., 2021), and YFCC-15M (Thomee et al., 2016); we use the parser implemented in nltk (Bird, 2006). At this time, the concept set size is $|C|\approx$ 20M. Then, we filter out nonessential concepts from the large base concept set according to policies based on Oikarinen et al. (2023); please see Appendix A. Finally, after filtering concepts, we obtain the concept bank containing $|C|\approx$ 5M concepts. We also discuss the effect of varying caption datasets used for collecting concepts in Sec. 5.5 and 5.6.2.

### v2

@@ Abstract
(a) Overview of Z-CBMs
(b) Comparison of Intervention
Figure 1: Zero-shot concept bottleneck models (Z-CBMs). (a) Z-CBMs predict concepts for input by retrieving them from a large-scale concept bank. Then, Z-CBMs predict labels based on the weighted sum of the retrieved concept vectors with importance weights yielded by sparse linear regression. (b) Z-CBMs can be intervened through modifying retrieved concepts and adding arbitrary concepts, whereas conventional CBMs allow only interventions in static concepts. The diverse and dynamic concepts produce accurate predictions and flexible collaborations with human experts.
@@ 1 Introduction
Developing human-interpretable models remains a primary interest within the deep learning research community. Concept bottleneck model (CBM, (Koh et al., 2020)) is an inherently interpretable neural network model, which aims to explain its final prediction via the intermediate concept predictions. Typically, CBMs are trained end-to-end on a target task to learn the input-to-concept and concept-to-class mappings. A concept consists of high-level semantic vocabulary for describing objects of interest in input data. For instance, CBMs can predict the final label “apple” from the linear combination of the concepts “red sphere,” ”green leaf,” and “glossy surface.” These intermediate concept predictions not only provide interpretability but also intervenability in the final prediction by editing the predicted concepts.
@@
In original CBMs (Koh et al., 2020), a concept set for each class label is defined by manual annotations, incurring massive labeling costs greater than ones of the class labels. To reduce these costs, Oikarinen et al. (Oikarinen et al., 2023) and Yuksekgonoul et al. (Yuksekgonul et al., 2023) automatically generate the concept sets by large language models (LLMs, e.g., GPT-3 (Brown et al., 2020b)) and use the multi-modal embedding space of vision-language models (VLMs, e.g., CLIP (Radford et al., 2021)) to learn the input-to-concept mapping through similarities in the multi-modal feature space. Although modern CBMs are free from manually pre-defined concepts, their practicality is still restricted by the requirements of training input-to-concept and concept-to-class mappings on target datasets. This means that CBMs have not been available without manually collecting target datasets and additional training of model parameters on them so far. Furthermore, CBMs allow interventions only in static concepts that are used in training, preventing human experts from flexible interactions with arbitrary concepts.
@@
To overcome these limitations, this paper introduces a novel problem setting of CBMs in a zero-shot manner for target tasks. In this setting, we can access pre-trained VLMs, but we cannot know the concepts composing target data in advance. This setting necessitates a two-stage zero-shot inference of input-to-concept and concept-to-class for unseen input samples. The zero-shot input-to-concept inference can not be solved by a naïve application of VLMs as the ordinary zero-shot classification of input-to-label, because it requires identifying a subset of relevant concepts from the large set of all concepts, rather than predicting a single label. Furthermore, the zero-shot concept-to-class inference is difficult because how to obtain the concept-to-class mapping is not obvious without target data and training, which are unavailable in this setting. Therefore, our primary research question is: How can we achieve interpretable and intervenable concept-based prediction through zero-shot input-to-concept and concept-to-class inference without relying on target datasets and additional training?
@@
We present a novel CBM class called zero-shot concept bottleneck models (Z-CBMs). Z-CBMs are zero-shot interpretable models that employ off-the-shelf pre-trained VLMs with frozen weights as the backbone (Fig. 1). Conceptually, Z-CBMs first perform concept retrieval to dynamically identify input-related concepts from a broad concept bank and then concept regression predicts the final label by simulating zero-shot classification capabilities of black-box VLMs via reconstructing the original input embedding from the retrieved concept embeddings. Our primary contribution is to achieve zero-shot input-to-concept and concept-to-class inference with this framework without additional training. Furthermore, our Z-CBMs allow interventions by arbitrary concepts described in natural language through the VLM feature spaces.
@@
We implement the components of Z-CBMs with simple yet carefully designed and effective techniques. For concept retrieval, Z-CBMs should cover broad domains to provide sufficient concepts for unseen inputs. To cover broad concepts, we build a large-scale concept bank, which is composed of millions of vocabulary extracted from large-scale text caption datasets such as YFCC (Thomee et al., 2016). Given an input sample, Z-CBMs dynamically retrieve concept candidates from the concept bank with an efficient and scalable cross-modal search algorithm. For concept regression, Z-CBMs estimate the importance of concepts for the input feature and then predict labels by the importance-weighted concept features. However, many of the retrieved concept candidates semantically overlap with each other, and thus, the semantically duplicated concepts with high importance by a naïve estimation method can harm the interpretability and intervenability for humans. To overcome this challenge, Z-CBMs find essential and mutually exclusive concepts for the final label prediction by leveraging sparse linear regression (e.g., lasso) to reconstruct the input visual feature vector by a weighted sum of the concept candidate vectors. Combining concept retrieval and concept regression enables Z-CBMs to predict final class labels with interpretable concepts for various domain inputs without any target training.
@@
Our extensive experiments on 12 datasets demonstrate that Z-CBMs can provide interpretable and intervenable concepts without any additional training. Specifically, we demonstrate that the sparse concepts identified by Z-CBMs exhibit strong correlation with input images and cover the annotated concepts in existing training-based CBMs. Furthermore, the Z-CBMs’ performance can be enhanced by human intervention in the predicted concepts, emphasizing the reliability of the concept-based prediction. We also show that Z-CBMs competitively perform with black box VLMs and existing CBMs with training. These results suggest the practicality of Z-CBMs for various domains.
2 Preliminaries
2.1 Concept Bottleneck Models (CBMs)
A CBM (Koh et al., 2020) is a classifier composed of a concept predictor $g:\mathcal{X}\to\mathcal{C}^{K}$ and a class label predictor $h:\mathcal{C}^{K}\to\mathcal{Y}$, where $\mathcal{X},\mathcal{C},\mathcal{Y}$ are input, concept, and class label spaces, and $K$ is the number of concepts. The goal is to predict the final class label $y\in\mathcal{Y}$ of input $x\in\mathcal{X}$ based on $K$ interpretable concepts $C=\{c_{i}\in\mathcal{C}\}_{i=1}^{K}$. To guarantee the interpretability and classification performance, $g$ and $h$ are jointly optimized on the following objective function (Koh et al., 2020):
$\min_{g,h}\mathbb{E}_{(x,C,y)\in\mathcal{D}}\left[\mathcal{L}(h\circ g(x),y)+\alpha\sum_{i}^{K}\mathcal{L}(g(x)_{i},c_{i})\right],$ (1)
where $\mathcal{D}$ is a training dataset, $\alpha$ is a hyperparameter, and $\mathcal{L}$ is a supervised loss function such as softmax cross-entropy loss. That is, CBMs’ interpretability is defined by their ability to detect concepts in the input accurately, which is obtained through the training of input-to-concept and concept-to-class predictions. In this sense, CBMs have the challenge of requiring human annotations of concept labels, which are more difficult to obtain than target class labels. Another challenge is potential performance degradation compared to backbone black-box models (Zarlenga et al., 2022; Moayeri et al., 2023; Xu et al., 2024) due to the difficulty of learning long-tailed concept distributions (Ramaswamy et al., 2023).
2.2 CBMs based on Vision-Language Models
To address the challenges, recent works (Yuksekgonul et al., 2023; Oikarinen et al., 2023; Yang et al., 2023) have focused on leveraging the capabilities of vision-language models (VLMs, e.g., CLIP (Radford et al., 2021)) and large language models (LLMs, e.g., GPT3 (Brown et al., 2020a)). These works automatically generate $C$ in text for each $(x,y)\in\mathcal{D}$ by prompting LLM, and then, train $g$ and $h$ using multi-modal feature spaces with a vision encoder $f_{\mathrm{V}}:\mathcal{X}\to\mathbb{R}^{d}$ and a text encoder $f_{\mathrm{T}}:\mathcal{T}\to\mathbb{R}^{d}$ provided by a VLM. We refer to such CBMs based on LLMs and VLMs as VLM-based CBMs. As pioneering works, Post-hoc CBMs (Yuksekgonul et al., 2023), Label-free CBMs (Oikarinen et al., 2023), and LaBo (Yang et al., 2023) firstly implemented this idea. The successor works have assumed the use of LLMs or VLMs, further advancing VLM-based CBMs (Panousis et al., 2023; Rao et al., 2024b; Tan et al., 2024; Srivastava et al., 2024). In particular, (Panousis et al., 2023) and (Rao et al., 2024a) are related to our work in terms of using sparse modeling to select concepts for input images. However, all of these existing VLM-based CBMs still require training specialized neural networks on target datasets, incurring additional training data and resources.
Handling the bi-level prediction in a zero-shot manner for unseen input is challenging because it can not be solved by naïve application of the existing zero-shot classification methods, which depend on limited vocabularies such as concepts related to ImageNet class names (Norouzi et al., 2014). Furthermore, current VLM-based CBMs and their variants (Bhalla et al., 2024) limit the number of concepts to a few thousand due to training and computational constraints, thereby restricting their generality.
In contrast to the previous VLM-based CBMs, the main purpose of this paper is to achieve fully zero-shot CBMs, which perform inference for input images from various domains without any additional training on target datasets.
@@ 3 Zero-shot Concept Bottleneck Models
In this section, we formalize the framework of Z-CBMs, which perform a zero-shot inference of input-to-concept and concept-to-class without target datasets and additional training (Fig. 1). Z-CBMs are composed of concept retrieval and concept regression. Concept retrieval finds a set of the most input-related concept candidates from millions of concepts by querying an input image feature with a semantic similarity search (Fig. 2). Concept regression estimates the importance scores of the concept candidates by sparse linear regression to reconstruct the input feature (Fig. 2). Finally, Z-CBMs provide the final label predicted by the reconstructed vector and concept explanations with importance scores.
@@
3.1 Zero-shot Inference Algorithm
Concept Retrieval. We first find the most semantically related concept candidates to input images from the large spaces in a concept bank $\mathbf{C}$ (Fig. 2). Given an input $x$, we retrieve the set of $K$ concept candidates $C_{x}\subset\mathbf{C}$ by using image and text encoders of pre-trained VLMs $f_{\mathrm{V}}$ and $f_{\mathrm{T}}$ as
$C_{x}=\underset{c\in\mathbf{C}}{\operatorname{Top-K}}~\operatorname{Sim}(f_{\mathrm{V}}(x),f_{\mathrm{T}}(c)),$ (2)
where $\operatorname{Top-K}$} is an operator yielding top-$K$ concepts in $\mathbf{C}$ from a list sorted in descending order according to a similarity metric $\operatorname{Sim}$. Throughout this paper, we use cosine similarity as $\operatorname{Sim}$ with normalized inputs by following (Conti et al., 2023). Thanks to the scalability of the similarity search algorithm (Johnson et al., 2019), Eq. (2) can efficiently find the concept candidates in an arbitrary concept bank $\mathbf{C}$, which contains millions of concepts to describe inputs in various domains.
Concept Regression. Given a concept candidate set $C_{x}=\{c_{1},...,c_{K}\}$, we predict the final label $\hat{y}$ by selecting essential concepts from $C_{x}$. Conventional CBMs infer the $C_{x}$-to-$\hat{y}$ mapping by training neural regression parameters on target tasks, which incurs the requirements of target dataset collections and additional training. Instead, we solve this task with a different approach, leveraging the zero-shot performance of VLMs. As shown in the previous studies (Radford et al., 2021; Jia et al., 2021), VLMs can be applied to zero-shot classification by inferring a label $\hat{y}$ by matching input $x$ and a class name text $t_{y}\in\mathcal{T}$ in the multi-modal feature spaces as follows.
@@ 3.1 Zero-shot Inference Algorithm
Through this objective, we can achieve $W$ not only for approximating image features but also for effectively estimating the contribution of each concept to the label prediction owing to the sparse regularization $\|W\|_{1}$. Since $C_{x}$ is retrieved from a large-scale concept bank $\mathbf{C}$, it often contains noisy concepts that are similar to each other, undermining interpretability due to semantic duplication. In this context, the sparse regularization enhances interpretability by penalizing and eliminating unimportant concepts for the label prediction (Hastie et al., 2015).
@@
[…]
@@
Table 1: Concept Accuracy on Bird (Welinder et al., 2010). Z-CBMs can infer large parts of ground-truth concepts without additional training.
Model Accuracy
CBM (Koh et al., 2020) 71.61
CDM (Panousis et al., 2023) 45.61
Z-CBM (Ours) 60.49
3.2 Feasibility Study
We show a preliminary experiment evaluating how much Algorithm 1 can accurately infer the ground-truth concepts, given the concept set from a fully annotated dataset as the concept bank $\mathbf{C}$. To this end, we tested CBM (Koh et al., 2020), CDM (Panousis et al., 2023) as a VLM-based CBM, and Z-CBM on the Bird dataset (Welinder et al., 2010), which has human-annotated concept labels. We used CLIP ViT-B/32 (Radford et al., 2021) as the backbone and the annotated 312 concepts as the concept bank $\mathbf{C}$ for CDM and Z-CBM. More detailed protocols are described in Appendix A. Table 1 shows the concept accuracy. Z-CBM outperformed CDM, which requires additional training, and achieved approximately 80% of the performance of CBM trained with the ground truth concepts. This indicates that Z-CBMs can find valid concepts without additional training by concept retrieval and regression. In practice, since full concept annotations are not generally given for unseen inputs, Z-CBMs adopt a large-scale concept bank that covers broad concepts for various domains.
@@ 4 Implementation
Vision-Language Models. Z-CBMs allow leveraging arbitrary pre-trained VLMs for $f_{\mathrm{V}}$ and $f_{\mathrm{T}}$. We basically use the official implementation of OpenAI CLIP (Radford et al., 2021) and the publicly available pre-trained weights. [^1] https://github.com/openai/CLIP Specifically, by default, we use ViT-B/32 as $f_{\mathrm{V}}$ and the base transformer with 63M parameters as $f_{\mathrm{T}}$ by following the original CLIP. In Section 5.6.1, we show that other VLM backbones (e.g., SigLIP (Zhai et al., 2023) and OpenCLIP (Cherti et al., 2023)) are also available for Z-CBMs.
Concept Bank Construction. Here, we introduce the construction protocols of the concept bank $C$ of Z-CBMs. Since Z-CBMs can not know concepts of input image features in advance, a concept bank should contain sufficient vocabulary to describe the various domain inputs. To this end, we extract concepts from multiple image caption datasets and integrate them into a single concept bank. Specifically, we automatically collect concepts as noun phrases by parsing each sentence in the caption datasets, including Flickr-30K (Young et al., 2014), CC-3M (Sharma et al., 2018), CC-12M (Changpinyo et al., 2021), and YFCC-15M (Thomee et al., 2016); we use the parser implemented in nltk (Bird, 2006). At this time, the concept set size is $|C|\approx$ 20M. Then, we filter out nonessential concepts from the large base concept set according to policies based on Oikarinen et al. (2023); please see Appendix A. Finally, after filtering concepts, we obtain the concept bank containing $|C|\approx$ 5M concepts. We also discuss the effect of varying caption datasets used for collecting concepts in Sec. 5.5 and 5.6.2.